1. Let A = {a,b,c}, B = {x,y}, and C = {0,1}. Find
  2. a) A×B×C. 
1
Expert's answer
2021-06-07T03:42:18-0400

Using the definition of the Cartesian product, we know that "A \times B \times C" are ordered triplets of any combination of the elements in each set (in the same order as the sets in the Cartesian product).

"A \times B \times C = \\
= {(a,x,0), (a,x,1), (a,y,0), (a,y,1), (b,x,0), (b,x,1), (b,y,0), (b,y,1), (c,x,0), (c,x,1), (c,y,0), (c,y,1)}"